# Nonlinear reduction strategies for data compression: a comprehensive comparison from diffusion to advection problems

### v1

@@ 3.2. Spectral Methods for Manifold Learning
Indeed, performing PCA on the feature space, we study the spectral decomposition of the matrix $\bm{C}:=\bm{\Phi}^{T}\bm{\Phi}$, where $\bm{\Phi}\in\mathbb{R}^{(n\times M)}$ is the evaluation matrix of the function $\Phi$ at the elements of the data manifold =$\mathcal{M}=\{u_{\mu_{i}}\}_{i=1}^{n}$. It can be verified that, being $\{\lambda^{\bm{C}}_{j}\}_{j=1}^{R}$ and $\{\bm{v}^{\bm{C}}_{j}\}_{j=1}^{R}$ respectively the eigenvalues and eigenvectors of the matrix $\bm{C}$ of rank $R$
@@
Multi-Dimensional Scaling $-\frac{1}{2}\bm{H}\bm{D}\bm{H}$
Isomap $-\frac{1}{2}\bm{H}\bm{D}^{(g)}\bm{H}$
@@ 3.2.1. Multi-Dimensional Scaling (MDS)
MDS [12] aims to preserve a notion of distance between points, which is encoded in the matrix $\bm{D}$ and then double-centered via the matrix $\bm{H}:=\bm{I}_{n}-\frac{1}{n}\bm{1}_{n}$. This is done by projecting the data on the directions which maximise the distance between the training points, similarly to what is done with the matrix $K$ in the classical case of PCA, to preserve the variance in the reduced representation.
@@ 3.2.2. Isomap
It follows the same principles as MDS, only changing the distance used to compute matrix $\bm{D}^{(g)}$, which is not the Euclidean distance, but the geodesic one, defined as the length of the shortest path between two points on the possibly curvy (i.e., nonlinear) manifold. It can be computed by approximating it with a sum of local Euclidean distances between neighbour points [5].
@@ 3.2.3. Spectral Clustering
The main idea behind Spectral Clustering (SC) is to divide the data $\bm{U}$ into clusters which are as separated as possible [41]. It is done through an optimization procedure based on the so-called adjacency matrix $\bm{W}$, defined as
@@ 3.2.4. Locally Linear Embedding
Among the techniques searching for a reduced representation of $\bm{U}$, the goal of Locally Linear Embedding (LLE) methods is to preserve the linear reconstruction weights of data neighbors [34]. Indeed such approach is divided in two fundamental optimization steps: the first in the high-dimensional space, and the second in the low-dimensional one. During the first stage, one searches for the weight matrix $\bm{\tilde{W}}$ defined as
@@
3.3. Auto-Encoders for Latent Space construction
@@
We can refer to the pipeline reported in Figure 1 also for their description. Indeed, they search for the reduced representation which minimize the reconstruction error on the training dataset, relying on the Universal Approximation Theorem (UAT) [18], that enables a representation through a deep network of a large variety of invertible encoding maps satisfying such minimization constraint. The fixed reduced dimension $N$ is expressed by construction by the number of neurons of the latent representation, and thus has to be established a priori to the training phase. In our notation, the encoder approximates the function $Z_{N}\equiv Z_{N,\mu}:\mathcal{X}\rightarrow\mathbb{R}^{N}$ that transforms the datum $u_{\mu_{i}}$ correspondent to the parameter $\mu_{i}$ in its latent representation $\bm{z}\in\mathbb{R}^{N}$, while the decoder corresponds to $Z_{N}^{-1}\equiv Z_{N,\mu}^{-1}$ that maps back into $\mathcal{X}$ the latent representation of the input. Nevertheless, in this case, we can not assume those maps to be linear with respect to $\mu$, and in general it is difficult to explicitly find the correspondent bases $\{\psi_{i}\}_{i=1}^{N}$ spanning the reduced space. Indeed, even assuming the encoder as a composition of a nonlinear map $\Phi$ and a simpler reduction step $Z_{N}$ and the decoder as the composition of the inverse functions, there is no easy way to characterize the space spanned by the reduced representation of the data.
@@
The second class instead aim to build an architecture resistant to input non-informative fluctuations, with the goal of inferring a stable latent representation. In order to do that, a regularizing term is inserted regarding the Jacobian matrix of the hidden layer $\bm{J}_{A}$, which represents the derivative of each hidden node with respect to each input component:
@@
Therefore, Auto-Encoders obtain a latent representation through loss minimizations that are not explicitly aware of the topological characteristics of the manifold, as instead happened in Section 3.2 with Manifold Learning techniques. On the other hand, such DL architecture focuses on enabling an accurate reconstruction of samples, which is not always possible in the previous cases. Furthermore, in many cases AutoEncoder provide an advantageous reduction architecture for Out Of Sample (OOS) estimation, i.e. for the identification of the reduced representation for data not belonging to the training set. Indeed, in all the approaches presented in Sections 3.1 and 3.2, one has to perform some sort of multivariate regression strategy to generalize the set of latent data $\{\mu_{i},\bm{z}_{i}\}_{i=1}^{N}$ to the actual reduction map $Z_{N}:u_{\mu}\mapsto\bm{z}_{\mu}\ \forall\mu\in\mathcal{P}$. In the DL case, instead, one can simply add additional layers to generalize this map, thus optimizing directly in the training phase also the reconstruction of the latent dynamics for generic values of $\mu$ [28].
@@ 4. Numerical Results
In this section, we apply the nonlinear reduction methods described above to the Advection-Diffusion problem presented in Section 2. In particular, we test them on the pure Advection problem ($c_{D}=0,c_{T}=4$), pure Diffusive problem ($c_{D}=0.1,c_{T}=0$), and Advection-Diffusion one ($c_{D}=0.1,c_{T}=4$). For simplicity, in each test case, the parameters are kept fixed and the time is evolved in the range $[0,T]$ with $T=0.5s$, and we fix $\Omega:=[-1,3]$. The training set is composed by $20$ solution samples $u_{t_{i}}$ for $t_{i}$ uniformly distributed in $[0,T]$, while the testing set by $200$ samples uniformly distributed in $[0,T]$.
@@
Figure 5. Reconstruction of train and test data with the correspondent $l_{2}$ error of PCA, Registration method, and AutoEncoder for pure Advection problem with $N=2$.
Figure 6. Reconstruction of train and test data with the correspondent $l_{2}$ error of PCA, Registration method, and AutoEncoder for pure Diffusion problem with $N=2$
Figure 7. Reconstruction of train and test data with the correspondent $l_{2}$ error of PCA, Registration method, and AutoEncoder for Advection-Diffusion problem with $N=2$
@@
Furthermore, in Figures 5, 6 and 7, we report the reconstruction with the associated error for the training and testing data, respectively for PCA, Registration methods, and “vanilla” AutoEncoder for small latent dimension ($N=2$). We construct our “vanilla” AutoEncoder with Linear Layers interleaved by nonlinear activation function ($\tanh$) for both encoder and decoder, made up also by an hidden layer of dimension $\frac{D}{2}$. The test data is reconstructed by applying the inverse transformations of $\Phi$ and $Z_{N}$ to the reduced representation $z\in\mathcal{Z}_{N}$, while the test reduced variables are obtained by Kernel Ridge Regression (kRR) based on inverse multiquadratic RBFs [42]. We remark that, in such context, no Manifold Learning strategy is reported due to the fact that in general the map $\Phi$ is not explicitly known, making it impossible to reconstruct the original sample $u_{\mu_{i}}$ from its latent representation $z_{\mu_{i}}$. The results highlight the potential of Registration Methods, which, with very few latent variables, provide good approximations in all cases, also comparing to AutoEncoders, while proposing an interpretable reduction pipeline.
@@ 5. Conclusions
In this work, we offer a panoramic view of the main ideas concerning nonlinear reduction. In order to do this, we introduce technical insights on various methods belonging to different research fields: Numerical Analysis, Metric Learning, and Representation Learning.
@@
Finally, we apply the described methods to the class of Advection-Diffusion problems and analyze their approximation capabilities. In particular, we conduct the analysis on three test cases: pure Diffusion, pure Advection, and Advection-Diffusion problems. We report the reconstruction errors for the classes of Nonlinear ROMs in Numerical Analysis (Registration Methods) and Representation Learning (AutoEncoders), and we analyze the spectral decay for metric learning approaches (kPCA). By analyzing the whole set of approaches, we observed that, as expected, while Diffusion problems seem to be naturally well-suited for linear reduction, Advection problems are successfully reduced only when a proper pre-processing stage is applied to the original manifold $\mathcal{M}$. In particular, LLE and Registration Methods seem to lead the way, suggesting to search for a $\Phi$ which transforms all the samples in a reference one or, more in general, that embeds some notion of distance between the training samples.

### v2

@@ 3.2. Techniques from Manifold Learning: Kernel and Spectral methods
Indeed, performing PCA on the feature space, we study the spectral decomposition of the matrix $\bm{C}:=\bm{\Phi}^{T}\bm{\Phi}$, where $\bm{\Phi}\in\mathbb{R}^{(n\times M)}$ is the evaluation matrix of the function $\Phi$ at the elements of the data manifold $\mathcal{M}=\{u_{\mu_{i}}\}_{i=1}^{n}$. It can be verified that, being $\{\lambda^{\bm{C}}_{j}\}_{j=1}^{R}$ and $\{\bm{v}^{\bm{C}}_{j}\}_{j=1}^{R}$ respectively the eigenvalues and eigenvectors of the matrix $\bm{C}$ of rank $R$
@@
Multi-Dimensional Scaling $\frac{1}{2}\bm{H}\bm{D}\bm{H}$
Isomap $\frac{1}{2}\bm{H}\bm{D}^{(g)}\bm{H}$
@@ 3.2.1. Multi-Dimensional Scaling (MDS)
MDS [17, 44] aims to preserve a notion of distance between points, which is encoded in the matrix $\bm{D}$ and then double-centered via the matrix $\bm{H}:=\bm{I}_{n}-\frac{1}{n}\bm{1}_{n}$. This is done by projecting the data on the directions which maximise the distance between the training points, similarly to what is done with the matrix $K$ in the classical case of PCA, to preserve the variance in the reduced representation.
@@ 3.2.2. Isomap
It follows the same principles as MDS, only changing the distance used to compute matrix $\bm{D}^{(g)}$, which is not the Euclidean distance, but the geodesic one, defined as the length of the shortest path between two points on the possibly curvy (i.e., nonlinear) manifold. It can be computed by approximating it with a sum of local Euclidean distances between neighbour points [9, 52, 4].
@@ 3.2.3. Spectral Clustering
The main idea behind Spectral Clustering (SC) is to divide the data $\bm{U}$ into clusters which are as separated as possible [55, 33]. It is done through an optimization procedure based on the so-called adjacency matrix $\bm{W}$, defined as
@@ 3.2.4. Locally Linear Embedding
Among the techniques searching for a reduced representation of $\bm{U}$, the goal of Locally Linear Embedding (LLE) methods is to preserve the linear reconstruction weights of data neighbors [45, 18, 57]. Indeed such approach is divided in two fundamental optimization steps: the first in the high-dimensional space, and the second in the low-dimensional one. During the first stage, one searches for the weight matrix $\bm{\tilde{W}}$ defined as
@@
3.3. Techniques from Representation Learning: Auto-Encoders
@@
We can refer to the pipeline reported in Figure 1 also for their description. Indeed, they search for the reduced representation which minimize the reconstruction error on the training dataset, relying on the UAT, that enables a representation through a deep network of a large variety of invertible encoding maps satisfying such minimization constraint. The fixed reduced dimension $N$ is expressed by construction by the number of neurons of the latent representation, and thus has to be established a priori to the training phase. In our notation, the encoder $f_{e}$ approximates the function $Z_{N}:\mathcal{X}\rightarrow\mathbb{R}^{N}$ that transforms the datum $u_{\mu_{i}}$ correspondent to the parameter $\mu_{i}$ in its latent representation $\bm{z}_{\mu_{i}}\in\mathbb{R}^{N}$, while the decoder $f_{d}$ corresponds to $Z_{N}^{-1}$ that maps back into $\mathcal{X}$ the latent representation of the input. Nevertheless, in contrast to the Registration and Manifold Learning cases, for Auto-Encoders we can not assume $\mathcal{Z}_{N}$ to be a linear space and $Z_{N}$ to be a projection operator with respect to some bases $\{\psi_{j}\}_{j=1}^{N}$, as it is nontrivial to explicitly identify such bases hypothetically spanning the latent space $Z_{N}$. Indeed, even assuming the encoder as a composition of a nonlinear map $\Phi$ and a reduction step $Z_{N}$, there is no easy way to characterize the latent space [28].
@@
The second class instead aims to build an architecture resistant to input non-informative fluctuations, with the goal of inferring a stable latent representation. In order to do that, a regularizing term is inserted regarding the Jacobian matrix of the hidden layer $\bm{J}_{A}$, which represents the derivative of each hidden node with respect to each input component:
@@
Therefore, Auto-Encoders obtain a latent representation through loss minimizations that are not explicitly aware of the topological characteristics of the manifold, as instead happened in Section 3.2 with Manifold Learning techniques. On the other hand, such DL architecture focuses on enabling an accurate reconstruction of samples, which, in contrast, is not always possible in the previous cases due to the lack of an explicit $\Phi^{-1}$ operator.
Furthermore, in many cases Auto-Encoders provide an advantageous reduction architecture for Out Of Sample (OOS) estimation, i.e. the identification of the reduced representation of data not belonging to the training set. Indeed, for all the approaches presented in Sections 3.1 and 3.2, one has to perform some sort of multivariate regression strategy on the set of “train reduced variables” $\{\mu_{i},\bm{z}_{i}\}_{i=1}^{n}$ to obtain the reduction operator $Z_{N}:u_{\mu}\mapsto\bm{z}_{\mu}\ \forall\mu\in\mathcal{P}$. In the DL case, instead, $Z_{N}$ is the encoder itself, which automatically maps any new input in a new point of the learnt reduced space.
@@ 4. Numerical Results
In this section, we apply the nonlinear reduction methods described above to the Advection-Diffusion problem described in Equation (2) (see Section 2). In particular, we test them on the pure Advection problem ($c_{D}=0,c_{T}=4$), pure Diffusive problem ($c_{D}=0.1,c_{T}=0$), and Advection-Diffusion one ($c_{D}=0.1,c_{T}=4$). For simplicity, in each test case, the parameters are kept fixed and the time is evolved in the range $[0,T]$ with $T=0.5s$, and we fix $\Omega:=[-1,3]$. The training set is composed by $20$ solution samples $u_{t_{i}}$ for $t_{i}$ uniformly distributed in $[0,T]$, while the testing set by $200$ samples uniformly distributed in $[0,T]$  [^1] The Python codes with the implementation of the methods exploited for the numerical results will be made available upon request..
@@
Figure 5. Reconstruction of train and test data shown in first and third row, respectively, with the correspondent $l_{2}$ errors (second and fourth row) of PCA (left-most column), Registration method (middle column), and AutoEncoder (right-most column) for pure Advection problem with $N=2$.
Figure 6. Reconstruction of train and test data shown in first and third row, respectively, with the correspondent $l_{2}$ errors (second and fourth row) of PCA (left-most column), Registration method (middle column), and AutoEncoder (right-most column) for pure Diffusion problem with $N=2$
Figure 7. Reconstruction of train and test data shown in first and third row, respectively, with the correspondent $l_{2}$ errors (second and fourth row) of PCA (left-most column), Registration method (middle column), and AutoEncoder (right-most column) for Advection-Diffusion problem with $N=2$
@@
Furthermore, in Figures 5, 6 and 7, we report the reconstruction with the associated error for the training and testing data, respectively for PCA, Registration methods, and “vanilla” AutoEncoder for small latent dimension ($N=2$). We construct our “vanilla” AutoEncoder with Linear Layers interleaved by nonlinear activation function ($\tanh$) for both encoder and decoder, made up also by an hidden layer of dimension $\frac{D}{2}$. In all the cases, the test data is reconstructed by applying the backward operators $\Phi^{-1}$ and $Z_{N}^{-1}$ to the “test reduced variables” $z_{\mu_{test}}\in\mathcal{Z}_{N}$, which are obtained through Kernel Ridge Regression (kRR) based on inverse multiquadratic RBFs [56] applied on the “train reduced variables” $\{\mu_{i},z_{i}\}_{i=1}^{n}$ collected during the offline phase. We remark that, in such context, no Manifold Learning strategy is reported due to the fact that in general the map $\Phi$ is not explicitly known, making it impossible to reconstruct the original sample $u_{\mu_{i}}$ from its latent representation $z_{\mu_{i}}$. The results highlight the potential of Registration Methods, which, with very few latent variables, provide good approximations in all cases, also comparing to AutoEncoders, while proposing an interpretable reduction pipeline.
@@ 5. Conclusions
In this work, we offer a panoramic view of the main ideas concerning nonlinear reduction. In order to do this, we introduce technical insights on various methods belonging to different research fields: Numerical Analysis, Manifold Learning, and Representation Learning.
@@
Finally, we apply the described methods to the class of Advection-Diffusion problems and analyze their approximation capabilities. In particular, we conduct the analysis on three test cases: pure Diffusion, pure Advection, and Advection-Diffusion problems. We report the reconstruction errors for the classes of Nonlinear ROMs in Numerical Analysis (Registration Methods) and Representation Learning (AutoEncoders), and we analyze the spectral decay for Manifold Learning approaches (kPCA). By analyzing the whole set of approaches, we observed that, as expected, while Diffusion problems seem to be naturally well-suited for linear reduction, Advection problems are successfully reduced only when a proper pre-processing stage is applied to the original manifold $\mathcal{M}$. In particular, LLE and Registration Methods seem to lead the way, suggesting to search for a $\Phi$ which transforms all the samples in a reference one or, more in general, that embeds some notion of distance between the training samples.